OpenStudy (anonymous):

Suppose you have 30 CDs. You know that you have 11 more CDs than your friend. Write and solve an equation to find the number of CDs your friend has. A. n – 11 + 30 = 30; 19 B. 30 + n = 11; –19 C. 30 + 11 = n; 41 D. n + 11 = 30; 19

OpenStudy (lilg132):

no point answering him hes just rude to people who try and help

OpenStudy (anonymous):

am not being rude , i just forgot to give a you a medal am new to this okay i didnt know sorry

OpenStudy (lilg132):

calling me an idiot doesnt help tho does it?

OpenStudy (akshay_budhkar):

that idiot was for me lol

OpenStudy (saifoo.khan):

ROFl.

OpenStudy (anonymous):

no it doesnt

OpenStudy (lilg132):

whoever it was for theres no need for it especially to people trying to help

OpenStudy (anonymous):

okay sorry

OpenStudy (lilg132):

oh what do you know the medals are rolling in now when theres no answers being given :)

OpenStudy (anonymous):

ok

OpenStudy (lilg132):

i accept your apology now do you want to learn how to do this or just the answers

OpenStudy (anonymous):

learn

OpenStudy (akshay_budhkar):

:D

OpenStudy (akshay_budhkar):

ok so let the number of Cds your friend has be n you have 11 more so you have n+11 but you have 30 that is given so n+11=30 and you have n=19

OpenStudy (anonymous):

Anyways for an answer.... Let n be the amount of CDs your friend has. You have 30 CDs and you 11 more than "n" Since you have both 30 CDs and 11 more CDs than n, those two values must be equal: 30 = 11 + n Subtract 11 from both sides of the equation: 30-11 = 11 - 11 + n and you get: n = 19.

OpenStudy (lilg132):

call the amount of cds your friend has n we know you have 11 more cds then your friend and we know the total amount of cds you have is 30 so you would write n + 11 = 30 now re arrange this to work out what n is you would get n = 30 - 11 n = 19

OpenStudy (anonymous):

oh i get it thanks its that i dont have really good teachers in my school and i didnt really get it